Diffusion Morning Report 02/17/09

Ted Berg tberg at stanford.edu
Tue Feb 17 06:23:55 PST 2009


Good Morning All,
    Looking OK this morning. TEOS 2 high pressure. SiGe temp cal. Poly 
added 10 microns since Friday down for tube clean tomorrow. Nitride not 
pumping.. Ebeam looking for beam. Enough for now.

  	  	  	  	  	 
  	Date: 	2/17/09 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	up 	  	  	  	 
Tylan 6-POCl3 	up 	  	  	  	 
Tylan 7-TEOS2 	down 	high pressure 	27.9/30 	  	 
Tylan 8-SiGe 	yellow 	temp cal 	1.12/30 	  	 
Tylan 9-Poly 	up 	down tomorrow for tube clean 	33.02/30 	  	 
Tylan 10-Nitride 	yellow 	not pumping 	.22/30 	  	 
Tylan 11-LTO 	up 	  	25.63/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	up 	  	  	  	 
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	  	  	  	 
TEL Spa 	up 	  	  	  	 
Burn Box 	up 	  	   	  	 
Ebeam 	down 	tip changed looking for beam 	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	  	B on line 130 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20090217/884e2dec/attachment.html>


More information about the diffusion mailing list